If f(x,y,z)=(x2y2−2y3z2−4x3z2+xyz) , the gradient of f at (1,−2,−1) is
1
Expert's answer
2020-11-01T14:53:05-0500

"f(x,y,z) = x^2y^2 - 2y^3z^2 - 4 x^3z^2 + xyz"

"grad f = (\frac{\partial f}{\partial x},\frac{\partial f}{\partial y},\frac{\partial f}{\partial z})"

"\frac{\partial f}{\partial x}= 2xy^2 - 12 x^2 z^2 + yz\\
\frac{\partial f}{\partial y} = 2x^2y - 6y^2z^2 + xz\\
\frac{\partial f}{\partial z} = -4y^3z-8x^3z + xy"


"gradf = (2xy^2 - 12 x^2 z^2 + yz ,\\2x^2y - 6y^2z^2 + xz,\\ -4y^3z-8x^3z + xy )"


And then put "(1,-2,-1)" into "gradf"


"grad f (1,-2,-1) = (-2,-29,-25)"